given points (2,4) , (x,8) slope = -1/2
How do I slove for x? Please show all your work.
Do you cross multiply?

Answer by stanbon(75887)

given points (2,4) , (x,8) slope = -1/2
How do I solve for x?
------------
Find "x":
slope = (8-4)/(x-2) = -1/2
----
4/(x-2) = -1/2
----
Cross multiply:
x-2 = 8
x = 10
==============
cheers,
Stan H.